A car dealer wants to draw a Pie Graph representing the different types of cars he sold in a given month. He sold a total of 90 this month with 15 of those cars being convertables. How many degrees should be used to represent convertables in the Pie Graph? (You do not have to use the degree symbol in your answer.)
Answer by Theo(13342)

the whole pie is 360 degrees.

total number of cars sold is 90.

15 are convertibles.

15/90 = 1/6 * total number of cars sold.

1/6 * 360 = 60 degrees.

60 / 360 = 1/6 * total number of degrees in circle.

the ratio of the number of convertibles to total number of cars is 15/90 = 1/6.

the ratio of the number of degrees to the total number of degrees in a circle is 60 / 360 = 1/6.

the ratios are the same.

60 degrees out of 360 degrees represents 15 out of 90 cars.